Exercise Minimizing handling work in storage

Weekly demand of one spare part from a big maintenance storage is 100 pieces, and in average the maintenance spare part order size is 2 pieces per order.

The purchasing lot size is 1000 pieces and the deliveries arrive when needed.

The picking time for this type of items is 3 minutes per order row and handling time in incoming material 20 minutes per delivery.  

a.Calculate the yearly picking time and handling time of incoming deliveries for this item.

b.If this fast mowing item is placed far from active picking area, the picking time is easily 20 % longer than if it is located in the actual picking area.   Calculate the extra cost if these items are not located correctly, when you know that the salary for storage worker is € 10 an hour + 65 % side cost.
Answer #1

A.

no. of week in year = 52 weeks

weekly demand of 1 spare parts = 100

yearly demand = 100*52 =5200

maintnance spare parts required = 5200*2=10400

purchasing lot size =10400/1000=10.4, or 11

yearly picking time = 11*3 =33 minute

total handling time = 11*20= 220 minute

B.

this fast moving item is placed far from active picking area then,

extra picking time = 33minute *20% = 7 minute

extra cost of worker = 7/60*€10 +65% = €2